# Reshuffling Resampling Splits Can Improve Generalization of Hyperparameter Optimization

### v1

@@ E.1 Datasets
Note that datasets serve as data generating processes (DGPs, Hothorn et al., 2005). As we are mostly concerned with the actual generalization performance of the final best HPC found during HPO based on validation performance we rely on a comparably large held out test set that is not used during HPO. We therefore use $5000$ data points sampled from a DGP as an outer test set. To further be able to measure the generalization performance robustly for varying data sizes available during HPO, we construct concrete tasks based on the DGPs by sampling subsets of (train_valid; $n$) size $500$, $1000$ and $5000$ from the DGPs. This results in 30 tasks in total (10 DGPS $\times$ 3 train_valid sizes). For more details and the concrete implementation of this procedure, see Appendix E.3. We also collected another $5000$ data points as an external validation set, but did not use it. Therefore, we had to tighten the restriction to $10000$ data points mentioned in the main paper to $15000$ data points as the lower bound on data points. To allow for stronger variation over different repetitions, we decided to use $20000$ as the final lower bound
E.2 Learning Algorithms
@@
The Elastic Net is based on sklearn’s Logistic Regression Classifier. We train it for a maximum of 1000 iterations using the “saga” solver. Table 7 lists the search space used during HPO.
@@
E.3 Exact Implementation
In the following, we outline the exact implementation of performing one HPO run for a given learning algorithm on a concrete task (dataset $\times$ train_valid size) and a given resampling. We release all code to replicate benchmark results and reproduce our analyses via https://github.com/sumny/reshuffling. For a given replication (in total 10):
@@ E.3 Exact Implementation
    • The learning algorithm is trained on training splits and evaluated on validation splits according to the resampling strategy. In case reshuffling is turned on, the training and validation splits are recreated for every HPO. We compute the Accuracy, ROC AUC and logloss when using a random search and compute ROC AUC when using HEBO and average performance over all folds for resamplings involving multiple folds.
@@
    We evaluate 500 HPCs when using random search and 250 HPC when using HEBO.
@@
As mentioned above, the test set is only varied for each of the 10 replica (repetitions with different seeds), but consistent for different tasks (i.e. the different learning algorithms are evaluated on the same test set, similarly, also the different dataset subsets all share the same test set). This allows for fair comparisons of different resamplings on a concrete problem (i.e. a given dataset, train_valid size and learning algorithm). Additionally, for the random search, the 500 HPCs evaluated for a given learning algorithm are also fixed over different dataset and train_valid size combinations. This is done to allow for an isolation of the effect, the concrete resampling (and whether it is reshuffled or not) has on generalization performance, reducing noise arising due to different HPCs. Learning algorithms themselves are not explicitly seeded to allow for variation during model training over different repetitions. Resamplings and partitioning of data are always performed in a stratified manner with respect to the target variable.
@@
For running HEBO, each resampling (standard and reshuffled for holdout, 5-fold holdout, 5-fold CV, 5x 5-fold CV) has to be actually run due to the adaptive nature of BO.
@@
For the random search experiments, this results in 10 (DGPs) $\times$ 3 (train_valid sizes) $\times$ 4 (learning algorithms) $\times$ 2 (holdout or 5x 5-fold CV) $\times$ 2 (standard or reshuffled) $\times$ 10 (repetitions) = 4800 HPO runs, [^9] Note that we do not have to take the 3 different metrics into account because random search allows us to simulate runs for different metric post hoc. each involving the evaluation of 500 HPCs and each evaluation of an HPC involving between 2 (for holdout; due to retraining on train validation data) and 26 (for 5x 5-fold CV; due to retraining on train validation data) model fits. In summary, the random search experiments involve the evaluation of 2.4 Million HPCs with in total 33.6 Million model fits.
@@
Similarly, for the HEBO experiments, this results in 10 (DGPs) $\times$ 3 (train_valid sizes) $\times$ 4 (learning algorithms) $\times$ 4 (holdout, 5-fold CV, 5x 5-fold CV or 5-fold holdout) $\times$ 2 (standard or reshuffled) $\times$ 10 (repetitions) = 9600 HPO runs, [^10] Note that HEBO was only run for ROC AUC as the performance metric., each involving the evaluation of 250 HPCs and each evaluation of an HPC involving of 2 (for holdout; due to retraining on train validation data) 6 (for 5-fold CV or 5-fold holdout; due to retraining on train validation data) or 26 (for 5x 5-fold CV; due to retraining on train validation data) model fits. In summary, the HEBO experiments involve the evaluation of 2.4 Million HPCs with in total 24 Million model fits.
E.4 Compute Resources
@@
We estimate our total compute time for the random search and HEBO experiments to be roughly 8.45 CPU years. Benchmark experiments were run on an internal HPC cluster equipped with a mix of Intel Xeon E5-2670, Intel Xeon E5-2683 and Intel Xeon Gold 6330 instances. Jobs were scheduled to use a single CPU core and were allowed to use up to 16GB RAM. Total emissions are estimated to be an equivalent of roughly 4636.74 kg CO2.
Appendix F Additional Benchmark Results
F.1 Main Experiments
@@
In this section, we provide additional visualizations of the results of our benchmark experiments. Below, we give an overview of the different types of analyses and visualizations we provide. Normalized metrics, i.e., normalized validation or test performance refer to the measure being scaled to $[0,1]$ based on the empirical observed minimum and maximum values obtained on the raw results level (ADTM, see Wistuba et al., 2018). More concretely, for each scenario consisting of a learning algorithm that is run on a given task (dataset $\times$ train_valid size) given a certain performance metric, the performance values (validation or test) for all resamplings and optimizers are normalized on the repetition level to $[0,1]$ by subtracting the empirical best value and dividing by the range of performance values. Therefore a normalized performance value of 00 is best and $1$ is worst. Note that we additionally provide further aggregated results on the learning algorithm level and raw results of validation and test performance via https://github.com/sumny/reshuffling.
@@
    • Rank w.r.t test performance in Figure 9.
  • HEBO vs. random search holdout
@@
    • Rank w.r.t test performance in Figure 13.
  • HEBO vs. random search 5-fold holdout
@@
    • Rank w.r.t test performance in Figure 17.
  • HEBO vs. random search 5-fold CV
@@
    • Rank w.r.t test performance in Figure 21.
  • HEBO vs. random search 5x 5-fold CV
@@
    • Rank w.r.t test performance in Figure 25.
@@
Figure 6: Random search. Average normalized performance over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 7: Random search. Average normalized test performance over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 8: Random search. Average improvement (compared to standard 5-fold CV) with respect to test performance of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 9: Random search. Average ranks (lower is better) with respect to test performance over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 10: HEBO vs. random search for holdout. Average normalized validation performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 11: HEBO vs. random search for holdout. Average normalized test performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 12: HEBO vs. random search for holdout. Average improvement (compared to standard holdout) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 13: HEBO vs. random search for holdout. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 14: HEBO vs. random search for 5-fold holdout. Average normalized validation performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 15: HEBO vs. random search for 5-fold holdout. Average normalized test performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 16: HEBO vs. random search for 5-fold holdout. Average improvement (compared to standard 5-fold holdout) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 17: HEBO vs. random search for 5-fold holdout. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 18: HEBO vs. random search for 5-fold CV. Average normalized validation performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 19: HEBO vs. random search for 5-fold CV. Average normalized test performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 20: HEBO vs. random search for 5-fold CV. Average improvement (compared to standard 5-fold CV) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 21: HEBO vs. random search for 5-fold CV. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 22: HEBO vs. random search for 5x 5-fold CV. Average normalized validation performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 23: HEBO vs. random search for 5x 5-fold CV. Average normalized test performance (ROC AUC) over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 24: HEBO vs. random search for 5x 5-fold CV. Average improvement (compared to standard 5x 5-fold CV) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 25: HEBO vs. random search for 5x 5-fold CV. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
F.2 Ablation on M-fold holdout
@@
Based on the 5x 5-fold CV we results we further simulated different $M$-fold holdout resamplings (standard and reshuffled) by taking M repeats from the first fold of the 5x 5-fold CV. This allows us to get an understanding of the effect more folds have on $M$-fold holdout, especially in the context of reshuffling.
@@
Figure 26: Random search. Average normalized validation performance over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 27: Random search. Average normalized test performance over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 28: Random search. Average improvement (compared to standard 5-fold holdout) with respect to test performance of the incumbent over tasks, learners and repetitions for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.

### v2

@@ F.1 Datasets
Note that datasets serve as data generating processes (DGPs; Hothorn et al., 2005). As we are mostly concerned with the actual generalization performance of the final best HPC found during HPO based on validation performance we rely on a comparably large held out test set that is not used during HPO. We therefore use $5000$ data points sampled from a DGP as an outer test set. To further be able to measure the generalization performance robustly for varying data sizes available during HPO, we construct concrete tasks based on the DGPs by sampling subsets of (train_valid; $n$) size $500$, $1000$ and $5000$ from the DGPs. This results in 30 tasks in total (10 DGPS $\times$ 3 train_valid sizes). For more details and the concrete implementation of this procedure, see Appendix F.3. We also collected another $5000$ data points as an external validation set, but did not use it. Therefore, we had to tighten the restriction to $10000$ data points mentioned in the main paper to $15000$ data points as the lower bound on data points. To allow for stronger variation over different replications, we decided to use $20000$ as the final lower bound.
F.2 Learning Algorithms
@@
The Elastic Net is based on sklearn’s Logistic Regression Classifier. We train it for a maximum of 1000 iterations using the "saga" solver. Table 8 lists the search space used during HPO.
@@
F.3 Exact Implementation
In the following, we outline the exact implementation of performing one HPO run for a given learning algorithm on a concrete task (dataset $\times$ train_valid size) and a given resampling. We release all code to replicate benchmark results and reproduce our analyses via https://github.com/slds-lmu/paper_2024_reshuffling. For a given replication (in total 10):
@@ F.3 Exact Implementation
    • The learning algorithm is trained on training splits and evaluated on validation splits according to the resampling strategy. In case reshuffling is turned on, the training and validation splits are recreated for every HPO. We compute the Accuracy, ROC AUC and logloss when using a random search and compute ROC AUC when using HEBO or SMAC3 and average performance over all folds for resamplings involving multiple folds.
@@
    We evaluate 500 HPCs when using random search and 250 HPC when using HEBO or SMAC3 (SMAC4HPO facade).
@@
As mentioned above, the test set is only varied for each of the 10 replica (repetitions with different seeds), but consistent for different tasks (i.e. the different learning algorithms are evaluated on the same test set, similarly, also the different dataset subsets all share the same test set). This allows for fair comparisons of different resamplings on a concrete problem (i.e. a given dataset, train_valid size and learning algorithm). Additionally, for the random search, the 500 HPCs evaluated for a given learning algorithm are also fixed over different dataset and train_valid size combinations. This is done to allow for an isolation of the effect, the concrete resampling (and whether it is reshuffled or not) has on generalization performance, reducing noise arising due to different HPCs. Learning algorithms themselves are not explicitly seeded to allow for variation during model training over different replications. Resamplings and partitioning of data are always performed in a stratified manner with respect to the target variable.
@@
For running HEBO or SMAC3, each resampling (standard and reshuffled for holdout, 5-fold holdout, 5-fold CV, 5x 5-fold CV) has to be actually run due to the adaptive nature of BO.
@@
For the random search experiments, this results in 10 (DGPs) $\times$ 3 (train_valid sizes) $\times$ 4 (learning algorithms) $\times$ 2 (holdout or 5x 5-fold CV) $\times$ 2 (standard or reshuffled) $\times$ 10 (replications) = 4800 HPO runs, [^10] Note that we do not have to take the 3 different metrics into account because random search allows us to simulate runs for different metric post hoc. each involving the evaluation of 500 HPCs and each evaluation of an HPC involving either 2 (for holdout; due to retraining on train validation data) or 26 (for 5x 5-fold CV; due to retraining on train validation data) model fits. In summary, the random search experiments involve the evaluation of 2.4 Million HPCs with in total 33.6 Million model fits.
@@
Similarly, for the HEBO and SMAC3 experiments, this each results in 10 (DGPs) $\times$ 3 (train_valid sizes) $\times$ 4 (learning algorithms) $\times$ 4 (holdout, 5-fold CV, 5x 5-fold CV or 5-fold holdout) $\times$ 2 (standard or reshuffled) $\times$ 10 (replications) = 9600 HPO runs [^11] Note that HEBO and SMAC3 were only run for ROC AUC as the performance metric., each involving the evaluation of 250 HPCs and each evaluation of an HPC involving either 2 (for holdout; due to retraining on train validation data), 6 (for 5-fold CV or 5-fold holdout; due to retraining on train validation data) or 26 (for 5x 5-fold CV; due to retraining on train validation data) model fits. In summary, the HEBO and SMAC3 experiments each involve the evaluation of 2.4 Million HPCs with in total 24 Million model fits.
F.4 Compute Resources
@@
We estimate our total compute time for the random search, HEBO and SMAC3 experiments to be roughly 11.86 CPU years. Benchmark experiments were run on an internal HPC cluster equipped with a mix of Intel Xeon E5-2670, Intel Xeon E5-2683 and Intel Xeon Gold 6330 instances. Jobs were scheduled to use a single CPU core and were allowed to use up to 16GB RAM. Total emissions are estimated to be an equivalent of roughly 6508.67 kg CO2.
Appendix G Additional Benchmark Results Visualizations
G.1 Main Experiments
In this section, we provide additional visualizations of the results of our benchmark experiments.
Figure 6 illustrates the trade-off between the final number of model fits required by different resamplings and the final average normalized test performance (AUC ROC) after running random search for a budget of $500$ hyperparameter configurations. We can see that the reshuffled holdout on average comes close to the final test performance of the overall more expensive 5-fold CV.
Figure 6: Trade-off between the final number of model fits required by different resamplings and the final average normalized test performance (AUC ROC) after running random search for a budget of $500$ hyperparameter configurations. Averaged over different tasks, learning algorithms and replications separately for increasing $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@ G.1 Main Experiments
Below, we give an overview of the different types of additional analyses and visualizations we provide. Normalized metrics, i.e., normalized validation or test performance refer to the measure being scaled to $[0,1]$ based on the empirical observed minimum and maximum values obtained on the raw results level (ADTM; see Wistuba et al., 2018). More concretely, for each scenario consisting of a learning algorithm that is run on a given task (dataset $\times$ train_valid size) given a certain performance metric, the performance values (validation or test) for all resamplings and optimizers are normalized on the replication level to $[0,1]$ by subtracting the empirical best value and dividing by the range of performance values. Therefore a normalized performance value of 00 is best and $1$ is worst. Note that we additionally provide further aggregated results on the learning algorithm level and raw results of validation and test performance via https://github.com/slds-lmu/paper_2024_reshuffling.
@@
    • Rank w.r.t. test performance in Figure 10.
  • HEBO and SMAC3 vs. random search holdout
@@
    • Rank w.r.t. test performance in Figure 14.
  • HEBO and SMAC3 vs. random search 5-fold holdout
@@
    • Rank w.r.t. test performance in Figure 18.
  • HEBO and SMAC3 vs. random search 5-fold CV
@@
    • Rank w.r.t. test performance in Figure 22.
  • HEBO and SMAC3 vs. random search 5x 5-fold CV
@@
    • Rank w.r.t. test performance in Figure 26.
@@
Figure 7: Random search. Average normalized performance over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 8: Random search. Average normalized test performance over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 9: Random search. Average improvement (compared to standard 5-fold CV) with respect to test performance of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 10: Random search. Average ranks (lower is better) with respect to test performance over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 11: HEBO and SMAC3 vs. random search for holdout. Average normalized validation performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 12: HEBO and SMAC3 vs. random search for holdout. Average normalized test performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 13: HEBO and SMAC3 vs. random search for holdout. Average improvement (compared to standard holdout) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 14: HEBO and SMAC3 vs. random search for holdout. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 15: HEBO and SMAC3 vs. random search for 5-fold holdout. Average normalized validation performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 16: HEBO and SMAC3 vs. random search for 5-fold holdout. Average normalized test performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 17: HEBO and SMAC3 vs. random search for 5-fold holdout. Average improvement (compared to standard 5-fold holdout) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 18: HEBO and SMAC3 vs. random search for 5-fold holdout. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 19: HEBO and SMAC3 vs. random search for 5-fold CV. Average normalized validation performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 20: HEBO and SMAC3 vs. random search for 5-fold CV. Average normalized test performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 21: HEBO and SMAC3 vs. random search for 5-fold CV. Average improvement (compared to standard 5-fold CV) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 22: HEBO and SMAC3 vs. random search for 5-fold CV. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 23: HEBO and SMAC3 vs. random search for 5x 5-fold CV. Average normalized validation performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 24: HEBO and SMAC3 vs. random search for 5x 5-fold CV. Average normalized test performance (ROC AUC) over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 25: HEBO and SMAC3 vs. random search for 5x 5-fold CV. Average improvement (compared to standard 5x 5-fold CV) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 26: HEBO and SMAC3 vs. random search for 5x 5-fold CV. Average ranks (lower is better) with respect to test performance (ROC AUC) of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
G.2 Ablation on M-fold holdout
@@
Based on the 5x 5-fold CV results we further simulated different $M$-fold holdout resamplings (standard and reshuffled) by taking M repeats from the first fold of the 5x 5-fold CV. This allows us to get an understanding of the effect more folds have on $M$-fold holdout, especially in the context of reshuffling.
@@
Figure 27: Random search. Average normalized validation performance over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 28: Random search. Average normalized test performance over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.
@@
Figure 29: Random search. Average improvement (compared to standard 5-fold holdout) with respect to test performance of the incumbent over tasks, learners and replications for different $n$ (train-validation sizes, columns). Shaded areas represent standard errors.